Maria's father is 3 times as old as Maria. Six year's ago, her father's age was 5 times as Maria's age. What are their age's now?
Answer by macston(5194)

M=Maria's age; F=Father's age=3M
F-6=5(M-6) (Maria and Father 6 years ago) Substitute for F
3M-6=5(M-6)
3M-6=5M-30 Add 30 to each side
3M+24=5M Subtract 3M from each side
24=2M Divide each side by 2
12=M ANSWER Maria is 12 years old
Father's age=3M=36 years old ANSWER Father's age is 36 years
CHECK
(36-6)=5(12-6)
30=5(6)
30=30